Pxsze:

A cylinder has a height of 7 inches. Its volume is 2,659.58 cubic inches. What is the radius of the cylinder? Use ​𝜋 ≈ 3.14 and round your answer to the nearest hundredth.

JamesTDG:

Formula to calculate radius: r = √(V / π h) Where, r = Radius of Cylinder V = Volume of Cylinder h = Height of the Cylinde Taken from EasyCalculation.com, please do the work on your own and not use their calculator.

Pxsze:

@AZ Take over

AZ:

okay haha so do you know the formula for volume of a cylinder? \(\text{volume} = \pi r h\) Have you seen that before?

AZ:

They tell you that the height (which is 'h' in the formula) is 7 The volume is 2,659.58 and to replace \(\pi\) with 3.14 and then you just have to to solve for 'r'

Pxsze:

when i slove for r its a long number..

AZ:

What do you get?

AZ:

I made a mistake in the formula it should be \(\text{volume} = \pi r^2 h\)

AZ:

so if you plug in the numbers they gave you \( 2659.58 = (3.14)\times r^2 \times 7\) what do you get if you solve for 'r'

Pxsze:

17.28577928

AZ:

No, how did you get that? First things first, we want to get the \( r^2\) all by itself so then that way we can eventually find \( r\) But first, let's simplify it what is 3.14 * 7 = ?

Pxsze:

its 21.98

Yourboisus:

rrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrreeeeeeeeeeeeee i hate cylinders

AZ:

now divide that on both sides so that way we can get r^2 all by itself so what is 2659.58 / 21.98 = ?

Yourboisus:

121

Pxsze:

121

Yourboisus:

then devide that by 3.14?

AZ:

so now we have \( r^2 = 121\) do you know what r = ? you have to take the square root on both sides

Yourboisus:

11

AZ:

That's your answer!

Yourboisus:

thank you very much

Pxsze:

:D Thanks big brain

AZ:

You're welcome :D

Yourboisus:

:D btw the problem was for me i asked him for help so he posted it

AZ:

oh, glad I could help!! :)

Yourboisus:

thx